# Repeated Padding as Data Augmentation for Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation aims to provide users with personalized suggestions based on their historical interactions. When training sequential models, padding is a widely adopted technique for two main reasons: 1) The vast majority of models can only handle fixed-length sequences; 2) Batching-based training needs to ensure that the sequences in each batch have the same length. The special value 0 is usually used as the padding content, which does not contain the actual information and is ignored in the model calculations. This common-sense padding strategy leads us to a problem that has never been explored before: Can we fully utilize this idle input space by padding other content to further improve model performance and training efficiency?
In this paper, we propose a simple yet effective padding method called Repeated Padding (RepPad). Specifically, we use the original interaction sequences as the padding content and fill it to the padding positions during model training. This operation can be performed a finite number of times or repeated until the input sequences’ length reaches the maximum limit. Our RepPad can be viewed as a sequence-level data augmentation strategy. Unlike most existing works, our method contains no trainable parameters or hyperparameters and is a plug-and-play data augmentation operation. Extensive experiments on various categories of sequential models and five real-world datasets demonstrate the effectiveness and efficiency of our approach. The average recommendation performance improvement is up to 60.3% on GRU4Rec and 24.3% on SASRec. We also provide in-depth analysis and explanation of what makes RepPad effective from multiple perspectives. The source code will be released to ensure the reproducibility of our experiments.
Data Augmentation; Sequential Recommendation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Although numerous models with different architectures have been proposed, padding has never been explored as a necessary technique for training sequential models. The vast majority of sequence models can only support inputting fixed-length sequences to define the model architecture and process model calculations. In addition, during training, batch processing techniques require all sequences in each batch to be the same length. Therefore, the maximum sequence length is usually used as a specified hyperparameter before the model is trained. During training, interceptions are performed for long sequences, and padding is performed for short sequences. Due to user habits and the long-tail effect, only a few users have long interaction sequences, while most users have very short ones. When padding a sequence, the special value 0 is usually treated as content. It does not contain meaningful information and is not involved in model calculations. This universally observed convention and custom leads us to the question: Can we fully utilize this idle input space by padding other content to improve model performance and training efficiency further?
@@
Motivated by the question above, we propose a simple yet effective padding strategy called Repeated Padding (RepPad for short) in this paper. Unlike the traditional padding strategy, we iteratively pad the original whole interaction sequence. This operation can be performed a limited number of times or repeated until the sequence length reaches a preset maximum limit. To avoid the situation where the head of the sequence is used to predict the tail, we add a delimiter between each padding. We want to utilize these input spaces that are not involved in model training but occupy positions to improve model performance and training efficiency.
@@
From another perspective, RepPad is a sequence-level data augmentation method. The training of recommendation models relies heavily on large-scale labeled data. Data sparsity has been one of the major problems faced by recommender systems (Jing et al., 2023). Earlier works use random data augmentation methods to increase the size of the training dataset directly, such as Dropout (Tan et al., 2016a) and Slide Windows (Tang and Wang, 2018). Besides, some researchers utilize counterfactual thinking (Wang et al., 2021), bidirectional Transformer (Jiang et al., 2021; Liu et al., 2021b), or diffusion model (Liu et al., 2023) to augment sequential data. Inspired by recent developments of Contrastive Learning (CL) (Jing et al., 2023; Yu et al., 2023), many effective data augmentation operators have been proposed (Xie et al., 2022; Liu et al., 2021a; Dang et al., 2023b; Qin et al., 2023; Tian et al., 2023). These models typically leverage shared encoders to encode original and augmented sequences. The representations of the original sequences are used for the main recommendation task, and the representations of the augmented sequence are used as the input for contrastive learning. It optimizes encoders by maximizing the agreements between ‘positive’ pairs, which are two augmentations from one sequence (Liu et al., 2021a). These methods also use the special value 0 as the padding content when dealing with sequences whose length is less than a specified value. Also, adding additional sequence data to the original dataset makes the training process more time-consuming. Our RepPad utilizes the idle padding space to perform repetitive padding before the sequences are fed into the model. There is no additional increase in the number of sequences that need to be involved in training.
@@
Different from elaborate data augmentation operators or the methods that require training, our method requires no training process and introduces no additional hyper-parameters or trainable parameters. It is independent of the backbone recommendation network and is a data augmentation plugin. We implement RepPad on various representative recommendation models based on different architectures. Extensive experimental results demonstrate the effectiveness and efficiency of our method. It also achieves satisfactory improvements over existing data augmentation methods. Furthermore, we explore the performance of different variants of RepPad. To explore what makes RepPad effective, we provide in-depth analysis from the perspectives of loss convergence and gradient stability.
@@
  • We also provide analysis and insights of why RepPad is effective from the perspectives of loss convergence and gradient stability. To our knowledge, this is the first work to explore the padding strategy in the recommendation field.
@@ 2.1. Problem Formulation
Suppose we have user and item sets denoted by symbols $\mathcal{U}$ and $\mathcal{V}$, respectively. Each user $u\in\mathcal{U}$ is associated with a sequence of interacted items in chronological order $s_{u}=[v_{1},\ldots,v_{j},\ldots,v_{\left|s_{u}\right|}]$, where $v_{j}\in\mathcal{V}$ indicate the item that user $u$ has interacted with at time step $j$ and $\left|s_{u}\right|$ is the sequence length. The sequential recommendation task can be formulated as follows. Given the sequences of interacted items $s_{u}$, how to accurately predict the most possible item $v^{*}$ that user $u$ will interact with at time step $\left|s_{u}\right|+1$, given by:
@@ 2.2. Padding in Sequential Recommendation
Padding is a common technique when training sequential models. The Batch process requires that the sequences in each batch have the same length. In addition, common CNN- or Transformer-based models can only handle fixed-length sequences and require model initialization and operations based on fixed sequence lengths. So, before training a recommendation model, we usually specify a maximum sequence length $N$. For sequences with an original length of more than $N$, the most recent $N$ items are usually chosen because they represent the user’s recent preferences. As shown in Fig.1 (a) and (b), original sequences of length less than $N$ are padded with the special value 0 to reach the length $N$. The special value 0 contains no actual information and does not participate in model calculations. The above process can be formulated as follows:
@@ 2.3. Sequence Data Augmentation
To alleviate the data sparsity problem, data augmentation is often used to augment the original data to increase the number of samples involved in model training. Given an original sequence $s_{u}$, a data augmentation operation makes minor but appropriate changes to $s_{u}$ to generate a new sequence $s_{u}^{d}$. Both $s_{u}$ and $s_{u}^{d}$ are used for model training or self-supervised learning. For example, as shown in Fig.1 (c), for an original sequence $s_{u}=[v_{5},v_{2},v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$, we intercept a portion of the consecutive sequence to generate a new sequence data $s_{u}^{d}=[v_{6},v_{4},v_{1},v_{7},v_{9}]$. For original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, we inert an item $v_{4}$ between $v_{7}$ and $v_{9}$ to generate sequence data $s_{u}^{d}=[v_{3},v_{7},v_{4},v_{9}]$. In each epoch, each sequence will be augmented at least once, so the total number of sequences involved in model training will be at least 2$\left|\mathcal{U}\right|$. The original and augmented sequences also need to go through the Eq. (2) before being fed into the model.
@@ 2.4. Ours: Repeated Padding
As illustrated in Fig. 1 (b) and (c), a large amount of free space is filled with 0 regardless of whether data augmentation is performed. Also, due to the long-tail effect, the interaction sequences of the vast majority of users are usually short. Therefore, our idea is to fully utilize these free spaces further to improve the efficiency of model training and recommendation performance.
@@
In order to achieve this objective, we propose a simple but effective padding strategy called Repeated Padding (RepPad). Specifically, we utilize the original sequence as padding content instead of the special value 0. As shown in Fig. 1 (d), give the original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, and we pad the original sequence once on the left and get the padded sequence $s_{u}^{p}=[v_{1},v_{2},v_{3},v_{4},v_{1},v_{2},v_{3},v_{4}]$. This operation can be formulated as:
where $m$ is padding times and |||| indicates sequence concatenation. This operation can be performed only once or repeated several times until there is not enough space left to hold the original sequence. In other words, $m$ can be pre-specified or calculated based on the original sequence length $s_{u}$ and the maximum sequence length $N$. For example, when the maximum sequence length is 10, the remaining length of 2 is insufficient to accommodate the original sequence of length 4, so it still needs to go through Eq. (2). The final input sequence $s_{u}^{f}=[0,0,v_{1},v_{2},v_{3},v_{4},v_{1},v_{2},v_{3},v_{4}]$. In addition, given the original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, the padding operation can be performed up to 3 times, and the final sequence $s_{u}^{f}=[0,0,v_{3},v_{7},v_{9},v_{3},v_{7},v_{9},v_{3},v_{7},v_{9}]$.
@@
Methods Augmentation Type No Hyper-Parameters No Training Model Agnostic Total Training Data
@@
However, there is a problem with the above padding strategy: there may be cases where the end of the original sequence is used to predict the head. For example, when use $[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},\underline{v_{1}},v_{2},v_{3},v_{4}]$ as training data, there will be cases where $v_{1},v_{2},v_{3},v_{4}$ is used to predict $v_{1}$. To solve this problem, we add the special value 0 between each repeated padding to prevent this from happening. This operation can be formulated as:
@@
As shown in Fig. 1 (e), given original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, the final sequence after Eq. (4) and (2) is $[0,v_{1},v_{2},v_{3},v_{4},0,v_{1},v_{2},v_{3},v_{4}]$. We give the possible values of $m$ and their interpretation in Table 1.
Our method has no trainable parameters or hyper-parameters that need to be manually tuned. It is a model-agnostic plug-and-play data augmentation approach. In the Experiments section, we will explore and analyze several variants of RepPad, including different times of padding and whether to add the delimiter 0. The best-performing variant is randomly selected between padding one time and the maximum number of times. Our method is very easy to implement and we present its Python procedure in Algorithm 1. So far, we have introduced the basic idea of our RepPad. It is worth emphasizing that the repeated padding sequences still need to go through the Eq. (2) operation before they can be fed into the model:
@@
  # The leave-one-out strategy is adopted.
@@
      max_pad_num = int(max_len / (len(orig_seq[:-3] + 1))
      # Randomly select a padding count.
@@
      # Perform the Repeated Padding operation.
@@
  # Then go through Eq.(2) before being fed into the model.
2.5. Discussions
2.5.1. Comparison with Existing Methods.
The advantages of our approach are as follows: (i) No increase in data size: From Fig. 1, we can observe that our approach performs data augmentation without increasing the amount of data for training. While previous approaches can be viewed as extending the length of the dataset vertically, our approach utilizes the free space horizontally. For traditional data augmentation methods, the amount of data involved in training will be at least 2$\left|\mathcal{U}\right|$. For sequential recommendation models that leverage both data augmentation and contrastive learning, each sequence is augmented at least twice, and with the addition of the original sequence, the number of model sequences involved in training is at least 3$\left|\mathcal{U}\right|$. (ii) No Hyper-parameters: These existing augmentation methods or augmentation operators (Slide Window, Dropout, Crop, Insert, Mask, Substitute, Reorder, and so on.) typically contain at least one hyper-parameter for each that must be manually tuned based on the dataset. However, our approach does not involve any hyperparameters that need to be tuned repeatedly. (iii) No Training Cost and Model Agnostic: Some researchers propose to use bidirectional Transformers (Liu et al., 2021b; Jiang et al., 2021) or diffusion models (Liu et al., 2023) for data augmentation. These methods require training of data augmentation modules in addition to training of backbone recommendation networks and have limitations on the type of backbone network. In contrast, our method does not contain any parameters, does not require a training process, and can be adapted to most sequential models.
We present a summary comparison between PepPad and existing methods in Table 2. In addition to training free and parameters free, our method does not increase the total amount of data involved in training. It should be emphasized that our focus is on the sequence-level data augmentation, which is also the most mainstream practice in data augmentation in the field of sequence recommendation. Augmentation at the embedding (Qiu et al., 2022) and model (Hao et al., 2023) levels is not involved in our discussion.
@@ 2.5. Discussions
Our approach intends to maximize the use of idle input space. Intuitively, RepPad is more effective on short-sequence datasets and may be ineffective on long-sequence datasets. This is because for long-sequence datasets, most users have sequence lengths that close to or exceed the given maximum sequence length, and thus there will be no space left for padding. We will verify our conjecture in the experimental section.

### v2

# Repeated Padding for Sequential Recommendation

@@ Abstract.
Sequential recommendation aims to provide users with personalized suggestions based on their historical interactions. When training sequential models, padding is a widely adopted technique for two main reasons: 1) The vast majority of models can only handle fixed-length sequences; 2) Batch-based training needs to ensure that the sequences in each batch have the same length. The special value 0 is usually used as the padding content, which does not contain the actual information and is ignored in the model calculations. This common-sense padding strategy leads us to a problem that has never been explored in the recommendation field: Can we utilize this idle input space by padding other content to improve model performance and training efficiency further?
In this paper, we propose a simple yet effective padding method called Repeated Padding (RepPad). Specifically, we use the original interaction sequences as the padding content and fill it to the padding positions during model training. This operation can be performed a finite number of times or repeated until the input sequences’ length reaches the maximum limit. Our RepPad can be considered as a sequence-level data augmentation strategy. Unlike most existing works, our method contains no trainable parameters or hyperparameters and is a plug-and-play data augmentation operation. Extensive experiments on various categories of sequential models and five real-world datasets demonstrate the effectiveness and efficiency of our approach. The average recommendation performance improvement is up to 60.3% on GRU4Rec and 24.3% on SASRec. We also provide in-depth analysis and explanation of what makes RepPad effective from multiple perspectives. Our datasets and codes are available at https://github.com/KingGugu/RepPad.
Data Augmentation; Sequential Recommendation; Padding
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 18th ACM Conference on Recommender Systems, October 14–18, 2024, Bari, Italy; October 14–18,2024; Bari, Italyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Although numerous models with different architectures have been proposed, padding has never been explored as a necessary technique for training sequential models. The vast majority of sequence models can only support inputting fixed-length sequences to define the model architecture and process model calculations. In addition, during training, the Batch Processing techniques require all sequences in each batch to be the same length. Therefore, the maximum sequence length is usually used as a specified hyperparameter before the model training. During training, interceptions are performed for long sequences, and padding is performed for short sequences. Due to user habits and the long-tail effect, only a few users have long interaction sequences, while most users have very short ones (Jiang et al., 2021). When padding a sequence, the special value 0 is usually treated as content. It does not contain meaningful information and is not involved in model calculations. This universally observed convention and custom leads us to the question: Can we utilize this idle input space by padding other content to improve model performance and training efficiency further?
@@
Motivated by the question above, we propose a simple yet effective padding strategy called Repeated Padding (RepPad for short) in this paper. Unlike the traditional padding strategy, we iteratively pad the original whole interaction sequence. This operation can be performed a limited number of times or repeated until the sequence length reaches a preset maximum limit. To avoid the situation where the head of the sequence is used to predict the tail, we add a delimiter between each padding. We aim to utilize these input spaces that are not involved in model training but occupying positions to improve model performance and training efficiency.
@@
From another perspective, RepPad is a sequence-level data augmentation method. The training of recommendation models relies heavily on large-scale labeled data. Data sparsity has been one of the major problems faced by recommender systems (Jing et al., 2023). Earlier works use random data augmentation methods to increase the size of the training dataset directly, such as Dropout (Tan et al., 2016a) and Slide Windows (Tang and Wang, 2018). Besides, some researchers utilize counterfactual thinking (Wang et al., 2021), bidirectional Transformer (Jiang et al., 2021; Liu et al., 2021b), or diffusion model (Liu et al., 2023) to augment sequence data. Inspired by recent developments of Contrastive Learning (CL) (Jing et al., 2023; Yu et al., 2023; Zhao et al., 2023a, b), many effective data augmentation operators have been proposed (Xie et al., 2022; Liu et al., 2021a; Dang et al., 2023b; Qin et al., 2023; Tian et al., 2023). These models typically leverage shared encoders to encode original and augmented sequences. The representations of the original sequences are used for the main recommendation task, and the representations of the augmented sequence are used as the input for contrastive learning (Liu et al., 2021a). These methods also use the special value 0 as the padding content when dealing with sequences whose length is less than a specified value. However, adding additional sequence data to the original dataset makes the training process more time-consuming. Our RepPad utilizes the idle padding space to perform repeated padding before the sequences are fed into the model. There is no additional increase in the number of sequences that need to be involved in training.
@@
In the discussion section, we analyze the advantages of RepPad over existing methods and present possible drawbacks. Different from elaborate data augmentation operators or the methods that require training, our method requires no training process and introduces no additional hyper-parameters or trainable parameters. It is independent of the backbone network and is a data augmentation plugin. We implement RepPad on various representative models based on different architectures. Extensive experimental results demonstrate the effectiveness and efficiency of our method. It also achieves satisfactory improvements over existing data augmentation methods. Furthermore, we explore the performance of different variants of RepPad. To explore what makes RepPad effective, we provide in-depth analyses from the perspectives of loss convergence and gradient stability.
@@
  • We provide analysis and insights of why RepPad is effective from the perspectives of loss convergence and gradient stability. To our knowledge, this is the first work to explore the padding strategy in the recommendation field.
@@ 2.1. Problem Formulation
Suppose we have user and item sets denoted by symbols $\mathcal{U}$ and $\mathcal{V}$, respectively. Each user $u\in\mathcal{U}$ is associated with a sequence of interacted items in chronological order $s_{u}=[v_{1},\ldots,v_{j},\ldots,v_{\left|s_{u}\right|}]$, where $v_{j}\in\mathcal{V}$ indicate the item that user $u$ has interacted with at time step $j$ and $\left|s_{u}\right|$ is the sequence length. The sequential recommendation task can be formulated as follows. Given the sequences of interacted items $s_{u}$, our task is to accurately predict the most possible item $v^{*}$ that user $u$ will interact with at time step $\left|s_{u}\right|+1$, formulated as follows:
@@ 2.2. Padding in Sequential Recommendation
Padding is a common technique when training sequential models. The Batch Processing requires that the sequences in each batch have the same length. In addition, common CNN- or Transformer-based models can only handle fixed-length sequences and require model initialization and operations based on fixed sequence lengths. So, before training a recommendation model, we usually specify a maximum sequence length $N$. For sequences with an original length of more than $N$, the most recent $N$ items are usually chosen because they represent the user’s recent preferences. As shown in Figure1 (a) and (b), original sequences of length less than $N$ are padded with the special value 0 to reach the length $N$. The special value 0 contains no actual information and does not participate in model calculations. The above process can be formulated as follows:
@@ 2.3. Sequence Data Augmentation
To alleviate the data sparsity problem, data augmentation is often used to augment the original data to increase the number of sequence involved in model training. Given an original sequence $s_{u}$, a data augmentation operation makes minor but appropriate changes to $s_{u}$ to generate a new sequence $s_{u}^{d}$. Both $s_{u}$ and $s_{u}^{d}$ are used for model training or self-supervised learning. For example, as shown in Figure1 (c), for an original sequence $s_{u}=[v_{5},v_{2},v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$, we intercept a portion of the consecutive sequence to generate a new sequence data $s_{u}^{d}=[v_{6},v_{4},v_{1},v_{7},v_{9}]$. For original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, we inert an item $v_{4}$ between $v_{7}$ and $v_{9}$ to generate sequence data $s_{u}^{d}=[v_{3},v_{7},v_{4},v_{9}]$. In each epoch, each sequence will be augmented at least once, so the total number of sequences involved in model training will be at least 2$\left|\mathcal{U}\right|$. The original and augmented sequences also need to apply function Eq. (2) before being fed into the model.
@@ 2.4. Ours: Repeated Padding
As illustrated in Figure 1 (b) and (c), a large amount of free space is filled with 0 regardless of whether data augmentation is performed. Also, due to the long-tail effect, the interaction sequences of the vast majority of users are usually short. Therefore, our idea is to utilize these free spaces further to improve the efficiency of model training and recommendation performance.
@@
Methods Augmentation Type No Hyper-Parameters No Training Model Agnostic Total Training Sequences
@@
In order to achieve this objective, we propose a simple but effective padding strategy called Repeated Padding (RepPad). Specifically, we utilize the original sequence as padding content instead of the special value 0. As shown in Figure 1 (d), give the original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, and we pad the original sequence once on the left and get the padded sequence $s_{u}^{p}=[\underline{v_{1},v_{2},v_{3},v_{4}},v_{1},v_{2},v_{3},v_{4}]$. This operation can be formulated as:
where $m$ is padding times and |||| indicates sequence concatenation. This operation can be performed only once or repeated several times until there is not enough space left to hold the original sequence. In other words, $m$ can be pre-specified or calculated based on the original sequence length $s_{u}$ and the maximum sequence length $N$. For example, when the maximum sequence length is 10, the remaining length of 2 is insufficient to accommodate the original sequence of length 4, so it still needs to apply function Eq. (2). The final input sequence $s_{u}^{f}=[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},v_{1},v_{2},v_{3},v_{4}]$. In addition, given the original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, the padding operation can be performed up to 3 times, and the final sequence $s_{u}^{f}=[0,0,\underline{v_{3},v_{7},v_{9}},\underline{v_{3},v_{7},v_{9}},v_{3},v_{7},v_{9}]$.
However, there is a problem with the above padding strategy: there may be cases where the end of the original sequence is used to predict the head. Using $[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},\underline{v_{1}},v_{2},v_{3},v_{4}]$ as training data, there will be cases where $v_{1},v_{2},v_{3},v_{4}$ is used to predict $v_{1}$. To solve this problem, we add the special value 0 between each repeated padding to prevent this from happening. This operation can be formulated as:
@@
As shown in Figure 1 (e), given original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, the final sequence after Eq. (4) and (2) is $[0,v_{1},v_{2},v_{3},v_{4},0,v_{1},v_{2},v_{3},v_{4}]$. We give the possible values of $m$ and their interpretation in Table 1.
Our method has no trainable parameters or hyper-parameters that need to be manually tuned. It is a model-agnostic plug-and-play data augmentation approach. In the Experiments section, we will explore and analyze several variants of RepPad, including different times of padding and whether to add the delimiter 0. The best-performing variant is The best-performing variant is randomly selected in the range from one to the maximum number of times. Our method is very easy to implement, and we present its Python procedure in Algorithm 1. So far, we have introduced the basic idea of our RepPad. It is worth emphasizing that the repeated padding sequences still need to apply function Eq. (2) operation before they can be fed into the model:
@@
2.5. Discussions
2.5.1. Comparison with Existing Methods.
The advantages of our approach are as follows: (i) No increase in number of sequences: From Figure 1, we can observe that our approach performs data augmentation without increasing the amount of data for training. While previous approaches can be considered as extending the length of the dataset vertically, our approach utilizes the free space horizontally. For traditional data augmentation methods, the amount of data involved in training will be at least 2$\left|\mathcal{U}\right|$. For sequential recommendation models that leverage both data augmentation and contrastive learning, each sequence is augmented at least twice, and with the addition of the original sequence, the number of model sequences involved in training is at least 3$\left|\mathcal{U}\right|$. (ii) No Hyper-parameters: These existing augmentation methods or augmentation operators (Slide Window, Dropout, Crop, Insert, Mask, Substitute, Reorder, and so on.) typically contain at least one hyper-parameter for each that must be manually tuned based on the dataset. However, our approach does not involve any hyperparameters that need to be tuned carefully. (iii) No Training Cost and Model Agnostic: Some researchers propose to use bidirectional Transformers (Liu et al., 2021b; Jiang et al., 2021) or diffusion models (Liu et al., 2023) for data augmentation. These methods require training of data augmentation modules in addition to training of backbone recommendation networks and have limitations on the type of backbone network. In contrast, our method does not contain any parameters, does not require a training process, and can be adapted to most sequential models. We present a summary comparison between PepPad and existing methods in Table 2. It should be emphasized that our focus is on sequence-level data augmentation, which is also the most mainstream practice in data augmentation in the field of sequential recommendation. Augmentation at the embedding (Qiu et al., 2022) and model (Hao et al., 2023) levels are not involved in our discussion.
@@ 2.5. Discussions
# The leave-one-out strategy is adopted. The last item (origseq[:-1]) in each sequence is the test set, and the penultimate item (orig_seq[:-2]) is the validation set.
@@
max_pad_num = int(max_len / (len(orig_seq[:-3]) + 1))
# Randomly select a padding count m.
@@
# Perform the Repeated Padding operation to get s_u^p.
@@
# Then apply function Eq.(2) before being fed into the model.
@@
Our approach intends to maximize the use of idle input space. Intuitively, RepPad is more effective on short-sequence datasets and may be ineffective on long-sequence datasets. This is because for long-sequence datasets, most users have sequence lengths that close to or exceed the given maximum sequence length, and thus, there will be no space left for padding. We will verify our conjecture in the experimental section.